OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x^3 + 4x^2 - 9x - 36

OpenStudy (***[isuru]***):

hmmm........... x^3 + 4x^2 - 9x - 36 = ( x -3) [ x^2 + 7x + 12 ] = ( x -3 ) [ x^2 + 4x + 3x + 12 ] = ( x - 3 ) [ x ( x + 4) + 3(x + 4 ) ] = ( x -3 ) [ (x + 3) ( x +4) ] make them equal to zero.. x= 3 or x = -3 or x = -2 which means the zeros r 3 , -3 , 4